LBC/Gregorian Technical Collimation Updates - 18 October 2011 UT

Observer(s): D. Thompson,A.Rakich, K. Boutsia, also present: A.!DiPaola, F.Pedichini, M.Centrone
Telescope Operator: D. Huerta (LBTO)
Telescope Support: K.Newton (LBTO)
Instrument Support: J. Morris (LBTO)

Summary

  • Checked LBC orientation...off but inconsistent, will check in daytime
  • LBC RPC error, lost 25m of data

  • Implemented new SX Gregorian collimation models (dated 20111018)
  • Verified active optics stayed low at all EL, good!
  • Collected field aberration data (in /OldRepository/AGW_Data/20111018/field)
  • Collected pointing data
  • AOM2 needs better astigmatism model!

Details

01:00 LUCI Collimation models from 2011A (dated 20110908) made with 8800 data points were updated to include offsets for the AOM2 from data taken 20111017. The New models were copied to the usual place with names of SXPMLUCIFERCollimation.20111018.dat and SXSMLUCIFERCollimation.20111018.dat. These have not yet been copied to the operational models.

2:45 Initial lbcrangebal had pointing corrections of ~10 arcsec. Offset SX M1 -0.8mm global offset to correct focus
dofpia, /first, /X2
Red converged on iteration 2, blue in 3

lbcrangebal 025302 025259
SX: dX = -0.450 dY = 0.227 dRX = 4.863 dRY = 9.653
DX: dX = 0.269 dY = -0.270 dRX = -5.790 dRY = -5.769
Old Delta New
------- ------- -------
CE -30.577 0.608 -29.969
CA -120.498 -7.904 -128.403

WCS: Red offset 0.266 degrees, blue says offset by 1.0515 degrees. Red zeropoint updated by 0.27d. Check says off by 1.5 degrees!

Fast OB created to do 3x1.0s exposures (RB_VV)
Blue Brot Red Rrot
30623 2.89 30620 2.35
30702 2.65 30651 2.08
30741 2.74 30722 2.16
31949 xxxx 31947 xxxx
32028 3.04 32017 2.46
32107 2.57 32049 2.09

AdP updating the rotator zero points, restarting LBCs.
3:48 Take 3 more pics

034843 2.85 034840 -2.53 034922 2.59 034912 -2.55 035001 2.58 034945 xxxx

lbcrangebal on WT10_364 (near M31 target field)
dofpia at target field
04:19 Execute M31_AndXXV_400s_1.ob B=3.2, R=3.3pixel FWHM
(BLUE Channel) OB execution error:
(TRACKERS) RPC Client/Host communication failed!
5   2011/10/18   04:34:41.841760         OB execution error: none [src/lbc/actions.c:1155]
4   2011/10/18   04:34:41.721731         >>> preset/expose/save interrupted at expose time
3   2011/10/18   04:34:41.721668         trackers system encountered a problem
2   2011/10/18   04:34:41.721590   TRACKERS      timeout on CCD Controller [src/trackers /trackers.c:2171]
1   2011/10/18   04:34:41.721511   TRACKERS      reinitializing CCD Controller failed
0   2011/10/18   04:34:20.361508   TRACKERS      RPC Client/Host communication failed retval:5 [src/trackers/trackers.c:1076] 

Blue hung during the second of three exposures, red continued through completion of all 3 exposures. Turn Off. 25m lost

04:45 Switch to Gregorian LFBG

Implement the new collimation models:
cp SXPMLUCIFERCollimation.20111018.dat SXPMLUCIFERCollimation.dat
cp SXSMLUCIFERCollimation.20111018.dat SXSMLUCIFERCollimation.dat

Poor initial collimation, Andrew checking. Seeing ~0.5" on guider. DIMM says

For a range of EL
ACTIVE preset, allow to collimate Record active optics in last wfscimage

Make list of selected rangebal'ed wfs images (e.g. "wcs.lis") hsel @wfs.lis "AO1COLL0,AO1COLL1,AO1COLL2,AO1COLL3,AO1COLL4,AO2COLL0,AO2COLL1,AO2COLL2,AO2COLL3,AO2COLL4,TELEL,T301,T201,AMBTEMP,GLASTEMP,$I" yes > "hdrAO.dat"

WT10_364 /Repository/20111018/GCS/left_wfscimage000033.fits

Found error in SXPMLUCIFERCollimation.20111018.dat, fixed. Clear out active optics, re-initialize, etc...start over.

/Repository/20111018/GCS/left_wfscimage000090.fits ACT0331 27.7
/Repository/20111018/GCS/left_wfscimage000099.fits WT10_357 34.9
/Repository/20111018/GCS/left_wfscimage000082.fits WT10_344 36.4
/Repository/20111018/GCS/left_wfscimage000108.fits ACT0299 40.9
/Repository/20111018/GCS/left_wfscimage000073.fits WT10_329 45.6
/Repository/20111018/GCS/left_wfscimage000126.fits ACT0366 48.2
/Repository/20111018/GCS/left_wfscimage000114.fits ACT0395 48.7
/Repository/20111018/GCS/left_wfscimage000067.fits WT10_342 52.8
/Repository/20111018/GCS/left_wfscimage000120.fits WT10_355 58.9
/Repository/20111018/GCS/left_wfscimage000062.fits WT10_327 61.9
/Repository/20111018/GCS/left_wfscimage000138.fits ACT0297 67.0
/Repository/20111018/GCS/left_wfscimage000056.fits WT10_353 69.9
/Repository/20111018/GCS/left_wfscimage000133.fits WT10_327 74.4
/Repository/20111018/GCS/left_wfscimage000044.fits WT10_364 75.6
/Repository/20111018/GCS/left_wfscimage000156.fits WT10_327 79.4
/Repository/20111018/GCS/left_wfscimage000152.fits ACT0327 79.6
/Repository/20111018/GCS/left_wfscimage000160.fits ACT0327 81.0
/Repository/20111018/GCS/left_wfscimage000147.fits ACT0340 82.0
/Repository/20111018/GCS/left_wfscimage000142.fits WT10_340 85.6

07:45 run field_collect using BS9107_field.list
  • preset failed because no star was found
  • correct pointing

07:50 field_collect,BS9107_field.list
===> Error in gcs_read_info: file not found
       /Repository/GCSfiles/dailyimages//left_wfscimage000168.info
===> Error in fits_gcs_to_struct: Information file not found
        /Repository/GCSfiles/dailyimages//left_wfscimage000168.fits
===> Error in fits_gcs_to_struct
         => Could not process /Repository/GCSfiles/dailyimages//left_wfscimage000168.fits
  => Collimated in 1 iterations
  • preset again
  • preset failed again after the first star
result from IIF 1
    setting reval val to 1
===> Active Optics loop is not running
  • repeat preset: field_collect,START=2
  • failed again trying to go to GS=0.
  • going to a POINTING STAR for pointing correction

08:17 field_collect - with the original BS9107_field.list, because now we do see it manually.
  • preset cancelled again, even though the star was visible.
  • problems with the AGW - homing the stages
  • it is taking a long time to home the stages... presets are failing also from iraf

08:41 killing the AGW and restarting because it is taking too long to home.
  • preset failed again - guide probe can't move to the position
  • kill the GCS and restart
  • preset failed again - AGW problem again

08:48 restart GCS and then restart AGW
  • pointing correction

08:55 field_collect,START=2

09:58 it failesd at GS=12 because it did not find the *info file
===> Error in gcs_read_info: file not found
       /Repository/GCSfiles/dailyimages//left_wfscimage000250.info
===> Error in fits_gcs_to_struct: Information file not found
        /Repository/GCSfiles/dailyimages//left_wfscimage000250.fits
===> Error in fits_gcs_to_struct
         => Could not process /Repository/GCSfiles/dailyimages//left_wfscimage000250.fits
  => Collimated in 1 iterations
  • it failed on axis because it found a fainter star and not BS9107 - pointing correction

10:00 field_collect,START=12
  • preset failed
  • we restart the AGW

10:07 field_collect,START=12

10:29 preset failed because of the AGW again
  • homing the stages

10:34 field_collect,START=15

10:40 failed because of the AGw again
  • this time we reboot the machine

10:48 field_collect,START=15
  • wrong star -pointing correction

10:50 field_collect,START=15

11:03 End of field_collection - BS9107_field.log
GStar wfsc #
0 176-182
23 183-185
0 186-187
GStar wfsc #
0 188-192
28 193-195
0 196-197
36 198-200
0 201-202
33 203-205
0 206
8 207-209
0 210
19 211-213
0 214-216
42 217-219
0 220-221
39 222-224
0 225-226
43 227-229
0 230-232
44 235-237
0 238-240
14 242-244
0 245-246
GStar wfsc #
0 255-258
12 259-261
0 262
41 263-265
0 266-268
22 269-271
0 272-274
GStar wfsc #
0 283-284
24 285-287
0 288-289
20 290-292
0 293-294

11:11 Collect pointing data for LFBG using AGw
  • Unwrap AZ to value >360.0
  • Correct pointing and collimation at ACT0370
  • Change GCS guider binning to 1x1, restart GCS (see AIP_L.bin2x2.cfg for old values)
  • Pointing log started: /home/telescope/PointingLogs/20111018_SX.log
  • newpoint * gs=0 Tel="GUIDE" WL=0.63 for all pointing stars.
  • AGw guide image TIME-OBS is ~6.87 minutes slow! 11:52:45.56 vs 11:59:38

1 ACT0370 3186
1 ACT0359 3200
3 ACT0400 3226
4 WT10_329 3240
5 WT10_278 3259
6 WT10_193 3276
7 WT10_267 3289
9 WT10_146 3303
10 WT10_286 3312
12 WT10_322 3332
13 ACT0335 3367
14 ACT0290 3388
15 ACT0256 3400
17 WT10_216 3415
19 WT10_161 3437
20 ACT0301 3453
21 WT10_265 3475
22 ACT0227 3491
23 WT10_312 3510
24 ACT0331 3527
25 ACT0402 3540
27 WT10_409 3560
28 ACT0346 3573
29 WT10_383 3589
31 WT10_399 3607
32 ACT0410 3627
33 WT10_357 3644
35 WT10_366 3663
36 WT10_355 3681
38 WT10_299 3702
39 WT10_229 3713
41 ACT0159 3723

AOM2 needs a better astigmatism model!

20111019 New pointing model generated: LUCIFERSX.20111018.ptmod, 2.81 arcsec rms. This is now set as the default, tcsdiscussion notified of the changes.

Summary from Andrew the following day

Date: Wed, 19 Oct 2011 14:50:28 -0700
Subject: Status of various collimation models.
From: Andrew Rakich 

Hi, following a summary of where we stand at our various focal stations:

LBC:

The default collimation model should be my latest one. It is so far very
well behaved, but still has a probelm with the thermal focus terms. On
typical nights this will mean that some focus adjustment needs to be done at
the start of the night, but once adjusted in it works very well.

SX Gregorian:

The latest collimation model should be the default one now for all SX
Gregorian. Dave can supply the name.


New collimation model is remarkably stable so far over three nights and
about a 5 degree temperature range. Focus in particular has only very rarely
slightly exceeded 125 microns of m2 z, usually less than 70 microns.

A pointing model has been made for this collimation model and seems to
behave satisfactoraly. Binodal astigmatism has been measured, but the
estimated corrections were not tested due to AGw 3 problems. So the
corrections have not been permanently applied as yet.

At its current level binodal may cause image gradient and defocus problems
that would be apparent in good seeing on when the guide star is far off-axis
at an unlucky orientation. We have a correction to try if observers
complain.
Note that the SX astigmatism model for M2 is not good, so we have modified
GCS so that only focus and astigmatism are modified until the combination of
these has dropped below 1000 nm. This change has been applied to Luci, Mods
and Pisces (still called IRTC) focal stations.

DX Gregorian:

Data was obtained last night for doing much needed elevation corrections for
Z7 and Z8. Z4 was amazingly well behaved, at about 8 degrees, Z terms were
20-30 microns all the time I was watching it. I'll be supplying a new range
balanced elevation corrected (for coma) model soon.

-- DavidThompson - 18 Oct 2011
Topic revision: r8 - 19 Oct 2011, JohnHill
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback